Find the equation of the line passing through the point (3,1) and parallel to the line which joins (3,-2) and (6,4).
Answer by stanbon(75887)

Find the equation of the line passing through the point (3,1) and parallel to the line which joins (3,-2) and (6,4).
----
slope of the given line = (4--2)/(6-3) = 2
----
Equation of line with slope = 2 passing thru (3,1).
--
Form: y = mx + b
Substitute for y,x,m and solve for "b":
1 = 2*3+b
b = -5
Equation:
y = 2x-5
Cheers,
Stan H.